OpenStudy (anonymous):

the square root of 32 is what

OpenStudy (theviper):

\[\Large{\sqrt{32}=\sqrt{2\times2\times2\times2\times2}=2\times2 \sqrt{2}=4\sqrt{2}}\]

OpenStudy (theviper):

\[\Large{4\sqrt{2}=4\times1.414}\]

OpenStudy (theviper):

Understood??

OpenStudy (unklerhaukus):

if you want to check \[(4\sqrt 2)^2=4^2\sqrt2^2=16\times2\]

OpenStudy (anonymous):

yea i get it

OpenStudy (theviper):

that is nice :)